40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 2404     Entry time: Wed Aug 28 16:51:49 2019
Author: anchal 
Type: DailyProgress 
Category: NoiseBudget 
Subject: Latest noisebudget as of Aug 23, 2019 

Attached is the latest noise budget calculated using the new object-oriented script.

Noise Budget Model:

Parameter files and all noise contribution PSD data.

Note that since we used Penn et al. JOSA B 36, 4, pp C15-C21  values for shear and bulk loss angles of the coating and used Hong et al . PRD 87, 082001 (2013) calculation procedure, as pointed in CTN:2401, the estimated coating brownian noise contribution is more than previously calculated and dominates the total estimated noise from about 20 Hz to 20 kHz. Also, the PLL noise was switched with Moku Frequency Noise measured in CTN:2357

Differences in the actual experiment from the model:

The experiment configuration file is attached. Major differences are:

  • Model uses 7mW of incident power. 8.8 mW was incident on North Cavity and 7.1 mW was incident on the South Cavity
  • Model uses 0.7+/-0.05 mode matching. North cavity was at ~0.65 and South was at ~0.77
  • Model uses temperature of 305 +/- 1 K. Vacuum can temperature was 306.53 K and table temperature was 292.23 K.

Beatnote Measurement Data


  • We aren't doing particularly better or worse than Nov 2017!
  • In fact there is a bump at around 500 Hz which was previously reported in CTN:2183 by Andrew and Craig as a scatter feature. So need to look into this.
  • FSS.... FSS!
Attachment 1: ExpStateAug23_2019Measurement_28-08-2019_175138_mean.yml  654 Bytes  Uploaded Wed Aug 28 18:53:41 2019  | Hide | Hide all
C3:PSL-NCAV_REFL_POW: '3.10606'
... 9 more lines ...
Attachment 2: CTNnoiseBudget_20190823.pdf  108 kB  Uploaded Wed Aug 28 18:58:59 2019  | Hide | Hide all
ELOG V3.1.3-